OpenStudy (anonymous):

Write an equation of the line in whose slope is 3 and passes through the point (1,2).

hartnn (hartnn):

hello @Brianawhodat :) \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) the equation of line with slope m and point (x1,y1) will be \(\large y-y_1=m(x-x_1)\) in your question, m=3, x1 = 1, y1=2 just plug in values!

OpenStudy (anonymous):

Thank you so much

hartnn (hartnn):

welcome ^_^ if you want me to verify your answer, just post what you got, i'll verify :)

OpenStudy (rishabh.mission):

for more such problem http://www.regentsprep.org/Regents/math/geometry/GCG1/EqLines.htm

OpenStudy (anonymous):

Marcy has a total of100 dimes and quarters. If the total value of the coins is $14.05, how many dimes and quarters does she have?

hartnn (hartnn):

Let q equal the number of quarters and d equal the number of dimes. Since total coins = 100, q + d = 100 -----Equation (1) At 25¢ for each q, the value of the quarters is 25q. At 10¢ for each d, the value of the dimes is 10d. Together they add up to 1405¢ so, other equation will be 25q + 10d = 1405 ----Equation (2) got these 2 equations ?? can you solve them ?

OpenStudy (anonymous):

Yes. Thank you again

hartnn (hartnn):

welcome ^_^ don't forget to click on 'Best Response' tab on any reply of anyone who helps you :) have fun learning with openstudy :)